# RACS and SADL: Towards Robust SMR in the Wide-Area Network

### v2

@@ VI Experimental evaluation
We use Amazon EC2 virtual machines [51] of type t2.xlarge (4 virtual CPUs, 16 GB memory) for replicas and clients, for WAN experiments. For LAN experiments, we use instances of type c4.4xlarge (16 virtual CPUs, 30 GB memory) for replicas and clients. We use Ubuntu Linux 20.04.5 LTS [52].
@@
Figure 6: Adversarial Performance in the WAN with 5 replicas – note that Multi-Paxos and Raft lines overlap
@@ VI-A Asynchronous Performance
This experiment evaluates RACS and SADL-RACS under simulated network attack scenarios, similar to attacks used in  [24, 25]. Our simulated attacker increases the egress packet latency of a minority of replicas chosen at random, by 500ms, dynamically, in time epochs. This experiment runs in the WAN setting with 5 replicas and 5 clients. We depict the results in Figure 6.
@@
Figure 7: Throughput versus latency for WAN normal-case execution, comparing pipelined RACS and SADL-RACS to pipelined Multi-Paxos, and pipelined EPaxos, with 3, 5 and 11 replica ensembles.
@@ VI-B RACS WAN Normal Case Performance
Figure 8: Throughput versus latency for WAN normal-case execution, comparing SADL-RACS to pipelined Multi-Paxos and pipelined RACS using 17B, 73B and 265B command sizes, using 5 replicas
@@ VI-D RACS LAN Normal Case Performance
Figure 9: Throughput versus latency for LAN normal-case execution, comparing RACS to Multi-Paxos, EPaxos, and Raft using 5 replicas
@@
RACS vs {Multi-Paxos and Raft}: We first observe that RACS achieves a saturation throughput of 420k cmd/sec, under a median latency upper bound of 10ms, that is comparable to the saturation throughput of Multi-Paxos (450k) and Raft (440k). Under normal case executions, RACS, Multi-Paxos, and Raft have 1 round-trip latency, serialized through a leader, hence provide comparable performance.
RACS vs EPaxos: Second, we observe that Epaxos-commit outperforms RACS, both in terms of latency and throughput. The EPaxos-commit experiment employs a conflict rate of 2%[55] hence 98% of the time, commands are committed in one round trip, without serializing through a leader replica. In contrast, RACS builds a total order of commands, serialized using a leader-replica, hence naturally the RACS performance is bottlenecked by the leader replica’s capacity.
VII Conclusion
We presented RACS, a novel randomized consensus protocol that simultaneously achieves three key goals. First, RACS ensures robustness under adversarial network conditions by utilizing a fallback mode of operation, replacing the traditional view change used by existing leader-based protocols. Second, RACS achieves performance comparable to existing leader-based protocols, under normal-case synchronous network conditions. Third, RACS adopts a simple design inspired by Raft, enabling easy engineering and seamless integration in cloud environments. RACS is the first protocol to simultaneously achieve these three goals.
We experimentally demonstrated RACS’s robustness and performance guarantees. RACS is suitable for adoption in the industry due to its robustness, high performance, and simplicity.
@@ A-A Definition
elected-fallback block: We refer to an fallback block $B_{f}$ generated in view $v$ with level $2$ as an elected-fallback block, if the common-coin-flip($v$) returns the index of the proposer $p_{l}$ who generated $B_{f}$ in the view $v$ and if the $<$asynchronous-complete$>$ for $B_{f}$ exists in the first $n-f$ $<$asynchronous-complete$>$ messages received. An elected-fallback block is committed same as a synchronously-committed block.

### v3

@@ VI Experimental evaluation
For WAN experiments, we chose t2.xlarge EC2 instances [51] with 4 vCPUs and 16 GB memory. For LAN experiments, we use instances of type c4.4xlarge (16 virtual CPUs, 30 GB memory) for replicas and clients. Since CPU is often the bottleneck in LAN settings—where network latency is low and available bandwidth is high—it is standard to use machines with more CPUs for LAN experiments and we followed this established practice. To ensure a fair, apples-to-apples comparison, we ran Raft, Multi-Paxos, and EPaxos on the same testbed. We use Ubuntu Linux 20.04.5 LTS [52].
@@
As the initial leader replica of Multi-Paxos [13], Raft [14], and RACS, we selected the replica located in Tokyo AWS region. We tested with other initial leader replica locations (Sydney, Seoul, Osaka, and Singapore), however, observed no noticeable performance difference. In Multi-Paxos and Raft, once a view timeout occurs, the next leader is selected uniformly at random – whoever initiates and successfully completes the view-change sub-routine becomes the leader for the subsequent view. In RACS, we use a deterministic round-robin next leader scheme for each new view.
@@
Figure 6: Adversarial Performance in the WAN with 5 replicas.
@@ VI-A Asynchronous Performance
A methodology to comprehensively evaluate consensus protocol performance under asynchronous network conditions remains an open research problem [55]. The only published method for asynchronous evaluation appears in [25], which we follow in this paper.
This experiment evaluates RACS and SADL-RACS under simulated network attack scenarios. The attacker increases the egress packet latency of a randomly selected minority of replicas by 500ms in dynamic time epochs. The experiment runs in a WAN setting with 5 replicas and 5 clients. We measured throughput, latency, downtime, and time-to-recover, but due to space limits and to align with recent literature [39, 25], we report only throughput and latency. We depict the results in Figure 6.
@@
Figure 7: Throughput versus latency for WAN normal-case execution for 3–11 number of replicas.
@@ VI-B RACS WAN Normal Case Performance
EPaxos is suited for applications that require only a partial order, such as key-value stores where per-key ordering is sufficient. In contrast, RACS is designed for applications that need a total order of all commands.
@@
Figure 8: Throughput versus latency for WAN normal-case execution using 17B, 73B and 265B command sizes.
@@ VI-D RACS LAN Normal Case Performance
Figure 9: Throughput versus latency for LAN normal-case.
@@
RACS vs {Multi-Paxos and Raft}: We observe that RACS achieves a saturation throughput of 420k cmd/sec, under a median latency upper bound of 10ms, that is comparable to the saturation throughput of Multi-Paxos (450k) and Raft (440k). Under normal case executions, RACS, Multi-Paxos, and Raft have 1 round-trip latency, serialized through a leader, hence provide comparable performance.
VII Conclusion, Limitations and Future Work
We presented RACS, the first randomized consensus protocol that simultaneously achieves three key goals; (1) robustness under adversarial network conditions, (2) performance comparable to existing leader-based protocols, under synchronous network conditions, and (3) simple design inspired by Raft, enabling seamless integration.
SADL-RACS has limitations that require future work. SADL-RACS requires at least four message hops per commit, making it less suitable for low-latency, high-throughput applications. We propose a DAG-based redesign that merges dissemination and ordering as a plausible method to achieve both low latency and high throughput.
While SADL-RACS optimizes for performance, robustness, and simplicity, it does not address energy efficiency—its chain-based design commits blocks even at low request rates, leading to high energy use. Although energy-efficient cloud systems are well studied, consensus protocol energy use remains unexplored, opening future directions in defining energy metrics, analyzing existing protocols, and designing energy-efficient consensus protocols.
Finally, RACS assumes that a majority of replicas are connected, however, loses liveness under extreme network partitions with only one or two connected replicas. Extending RACS to support such network partitions is left for future work.
Acknowledgements.
This work was supported by (1) Mysten Labs and the Sui Foundation, (2) the Austrian Science Fund (FWF) through the SFB SpyCode project F8512-N and by the WWTF through the project 10.47379/ICT22045, (3) EPFL research grants, (4) AXA Research Fund, and (5) Etherium grant.
@@ A-A Definition
elected-fallback block: We refer to an fallback block $B_{f}$ generated in view $v$ with level $2$ as an elected-fallback block, if the common-coin-flip($v$) returns the index of the proposer $p_{l}$ who generated $B_{f}$ in the view $v$ and if the $<$asynchronous-complete$>$ for $B_{f}$ exists in the first $n-f$ $<$asynchronous-complete$>$ messages received.